Candle-light dinner D. Leela and class-mates ”Come in”, shouted Leela, when she heard the knock at the door. Nivedita came in, looking scared.

”Why, what happened, Nivu?” asked Leela.

”The power switched off just as I was leaving the lift!” she exclaimed. If I had come even a second later, I would have been stuck in the lift. And I’m afraid of the dark,” she confessed.

”I like it when there’s a power-cut,” said Maya. ”I like to eat and play by candle-light.”

”Oh, no, it’s so dark and gloomy,” said Nivu. ”I agree. I like to have all the lights on in the room,” said Leela. In fact, I like to have all the lights on in all the rooms of the house!”

”Have you ever heard of global warming?” murmured Nivu. ”It’s not global warming I have to worry about; it’s Mother!” shot back Leela. ”She would never allow it!”

By this time, Maya had found a candle and match-box. She lighted the candle and looked around for a surface to put it on. ”Careful, Maya,” said Nivu. You’ll hurt yourself.” She picked up a plate and set it on the table. Maya carefully stuck the candle on the plate, with some melted wax. Then she sat on a chair and looked deep into the candle-flame. ”I like candle-light” she repeated.

”Just imagine, years ago, when there was no electricity. People must have huddled around candles just as we are doing now,” said Maya. Leela looked unhappy at the thought. ”We don’t have to live like they did. I am glad we have electricity and tap water.”

”Ah, water!” said Maya. She got a glass of water and carefully poured it into the plate with the candle. ”What are you doing, Maya?” asked Nivu. Maya looked at the candle in the water. ”I am in a boat,” she said, ”and it’s night and we have lighted the candle. There’s water all around us...”

”You funny thing!” exclaimed Nivedita, ”Why, you seem lost in another world. Perhaps I can also close my eyes and believe that we are on a boat!”

Leela said, ”That sounds fun!” Maya said, ”Mother showed me a nice trick with candle and water the other day. Do you want to see it?” ”Sure,” said Nivu. ”Can I do it too?”

”It’s actually a puzzle,” said Maya, ”so you can certainly try it. Here’s the question.” Saying which she ran away.

”What happened, where are you going?” asked Nivu. At that moment, Maya came back to the room, with a coin clutched in her hand. ”I went to ask Mother for a coin,” she said. She put the coin into the water on the plate.

”The puzzle doesn’t qquite work, I’m afraid. At least, you have to have just the right amount of water and I am not sure this is right. Let me ask the puzzle any way: how can you take the coin out of the water with your bare hands, and still not get your fingers wet?”

Nivu thought for a moment. ”Is it a trick?” ”Of course it’s a trick! All of them are,” said Maya. ”But it’s not cheating, just a nice trick.” Nivu looked at Leela. ”Don’t ask me” said Leela. Mother showed me this two years ago. I didn’t know she had shown it to Maya too.”

Maya said, ”It was when I was at home with cold and fever. I was bored so she showed me a whole lot of things. She showed me another trick with a candle flame, but I promised not to do it without an adult.”

”I know that one!” shouted Leela. ”Hey, hold it!” said Nivedita. I’m still on the first puzzle and I can’t think of any idea. Don’t go to the next one yet!”

”Actually,” said Maya, ”it may not be easy to figure out. I think I’ll show you the trick, and then Leela can do the other one.”

”That’s a great idea,” said Nivu. Maya went and got a glass tumbler. She held it upside-down over the candle flame. Nivu shouted out in alarm, ”Stop! You’ll put out the candle and we’ll be in the dark!”

Maya looked sheepish. ”Sorry, I forgot you don’t like the dark. Mother showed it to me in broad daylight, so it didn’t occur to me that I need to put out the candle!”

She brought another candle and lit it, setting it on the table. She again picked up the